# Abdoul Kader Kaboré

### v3

@@ 6.1 LLM-Based Test Generation: Strengths and Limitations
Recent research has explored the potential of Large Language Models (LLMs) for automated unit test generation, primarily focusing on correctness and code coverage. Siddiq et al. (2024a) examined ChatGPT-generated tests, emphasizing compilability and correctness issues, but their study did not evaluate key aspects such as maintainability, readability, or long-term usability. Similarly, Tang et al. (2024) and Yang et al. (2024) investigated coverage in LLM-generated tests, concluding that while LLMs can generate syntactically correct tests, they often fail at execution due to hallucinated dependencies and lack fault-detection capabilities.
Prior work focuses on static correctness, while we provide a comprehensive evaluation of LLM-generated tests. Our study assesses readability, maintainability, runtime failures, fault detection via mutation testing, and test smells, setting new benchmarks in software testing research.
6.2 Prompt Engineering for Test Generation
The impact of prompt engineering on LLM-based test generation has been explored in recent research. Yuan et al. (2024) and Chen et al. (2024) examined the effects of different prompting techniques on test correctness, demonstrating that Few-Shot Learning (FSL) improves correctness but does not fully mitigate hallucinated dependencies. Similarly, Yang et al. (2024) introduced Chain-of-Thought (CoT) prompting, highlighting its ability to enhance logical reasoning in test generation. However, these studies primarily focused on static correctness evaluations and did not analyze runtime failures, maintainability, or test smells, limiting their applicability in real-world testing workflows.
We systematically compare five prompting strategies—ZSL, FSL, CoT, ToT, and GToT—by executing generated tests and analyzing runtime outcomes (Pass, Fail, Timeout). Our study evaluates their impact on fault detection, test maintainability, and practical reliability, bridging the gap between theoretical correctness and real-world testing.
6.3 Execution-Based Correctness and Fault Detection
Prior works evaluating LLM-generated tests, such as Schafer et al.(Schäfer et al., 2023), Siddiq et al.(Siddiq et al., 2024a), and Chen et al. (Chen et al., 2024), have predominantly focused on static correctness criteria, including compilability and structural coverage. However, Palomba et al. (2016b) emphasized that high coverage does not necessarily imply meaningful tests—a concern validated by Bhatia et al. (2024), who found that while LLM-generated tests often execute successfully, they frequently lack effective fault-detection capabilities. To address this, we perform execution-based analysis by classifying test outcomes into Passed, Failed, and Timeout categories. We also apply mutation testing (Pitest) to evaluate fault detection—a previously overlooked aspect. Our findings show many LLM-generated tests include shallow assertions, limiting their defect detection effectiveness.
By going beyond static correctness checks, our study provides a real-world evaluation of test suite effectiveness, offering insights into critical weaknesses in LLM-generated test logic and execution reliability.
@@ 6.4 Readability, Maintainability, and Test Smell Analysis
While several prior studies acknowledge readability or maintainability issues in LLM-generated tests, they remain limited in scope. Siddiq et al. (2024a) mention readability concerns without structured evaluation, Yang et al. (2024) focus primarily on correctness metrics omitting test smell analysis, and Tang et al. (2024) and Yuan et al. (2023) report code style metrics without investigating maintainability impacts across diverse datasets.
Our study presents the first comprehensive multi-dimensional evaluation of LLM-generated test maintainability. We systematically analyze test smells using TsDetect, evaluate readability with a deep learning model tailored to JUnit code (Scalabrino et al., 2018), and assess structural conformance (MSR) and extraction feasibility (CSR)—dimensions not addressed in previous evaluations.
By integrating these metrics, our work uncovers maintainability challenges and provides actionable insights for prompt design and post-processing. This holistic evaluation addresses structural usability, setting a new benchmark for evaluating the practical viability of LLM-generated test suites.
@@ 6.5 Advancing LLM-Based Unit Test Improvement
Recent studies have explored LLM-driven unit test improvement techniques, primarily focusing on iterative refinement. Gu et al. (2024) introduced TestArt, a co-evolutionary approach that iteratively refines LLM-generated tests through repair iterations, enhancing correctness over multiple steps. Similarly, Alshahwan et al. (2024) investigated LLM-driven test improvements at Meta, focusing on refining correctness but without addressing key aspects such as maintainability, readability, or test smells. These studies demonstrate progress in test improvement but remain limited to correctness-focused refinements.
In contrast, our work provides a multi-dimensional assessment that goes beyond correctness refinements. We systematically evaluate test smell detection, an overlooked issue in prior work, revealing structural weaknesses in LLM-generated tests. Additionally, we incorporate execution-based correctness analysis, ensuring that refined tests not only compile but also behave correctly at runtime. Moreover, we introduce readability and maintainability metrics, assessing whether refined tests are understandable and practically usable in real-world software projects.
Our findings indicate that while iterative refinement can enhance LLM-generated tests, it does not fully address fundamental readability and execution challenges. This underscores the need for hybrid approaches, combining Search-Based Software Testing (SBST) techniques with LLM-based test generation to achieve both correctness and maintainability improvements.
@@ 6.6 Key Takeaways and Research Gap
Prior works on LLM-based test generation primarily focus on correctness and coverage, leaving key aspects unexplored. Our study fills this gap by providing a comprehensive evaluation framework that goes beyond static correctness checks. We incorporate execution-based correctness assessment, analyzing runtime behaviors through pass/fail/timeout classification and failure patterns. Additionally, we assess readability and maintainability using deep learning-based readability scoring, aligning with human evaluation.
Unlike previous studies that explore only one or two prompting techniques, we systematically compare five strategies—ZSL, FSL, CoT, ToT, and GToT. We also integrate mutation testing, a missing component in prior research, to evaluate fault detection capabilities. Beyond correctness and structural coverage (line, instruction, method), we analyze test smells, code extraction reliability (MSR & CSR), compilation errors, and runtime failures, uncovering key challenges in dependency resolution and API usage.
A key differentiator is our diverse dataset selection: SF110 (academic), Defects4J (real-world bugs), and CMD (industrial codebase). This ensures our findings generalize across different software environments. By bridging correctness, execution reliability, maintainability, dataset diversity, and fault detection, our study sets a new benchmark for evaluating LLM-generated test suites in
@@ 7 Discussion of Findings and Implications
Our study reveals that reasoning-based prompting (CoT, GToT) improves test correctness and maintainability, yet execution reliability and fault detection remain significant challenges. Many tests compile but fail to verify program behavior, while test smells affect maintainability. Hybrid approaches integrating SBST and human oversight are essential for robustness. Though prompt engineering improves reliability, hallucination-induced errors persist, indicating the need for better contextual awareness in handling dependencies and encapsulation. Execution-based validation is crucial, as static correctness alone is insufficient for real-world use. These findings emphasize the need for adaptive test generation pipelines incorporating automated validation, dependency resolution, and structured refinement.
@@ 7.1 Implications for Researchers
Our findings emphasize reasoning-based prompting’s significant role in improving LLM-generated test suites. Unlike prior studies (Siddiq et al., 2024a; Tang et al., 2024; Yang et al., 2024), which didn’t extensively analyze prompt engineering, we systematically evaluate ZSL, FSL, CoT, ToT, and GToT. Results show structured reasoning significantly enhances test adherence, compilability, and reliability, with GToT consistently outperforming other strategies.
Our multi-level coverage analysis reveals LLM-generated tests achieve moderate coverage but remain inferior to EvoSuite. This demonstrates the need for hybrid approaches integrating SBST techniques with LLMs. Our findings also challenge the assumption that larger models produce better tests, as GPT-3.5-Turbo occasionally outperforms GPT-4 on CMD, highlighting the importance of investigating how different architectures perform across diverse scenarios.
A critical takeaway is the oracle problem in test failure interpretation. While prior work focused on defect detection through failing tests, our analysis shows many ”Passed” tests are ineffective for fault detection, often targeting interfaces, abstract classes, or trivial assertions. This underscores the necessity of developing better validation techniques for meaningful defect identification.
Our readability and maintainability analysis reveals that while LLMs generate syntactically correct and readable tests, readability alone doesn’t ensure maintainability. Test smells—particularly Magic Number Tests and Assertion Roulette—persist across all models and prompting strategies, increasing debugging complexity. This highlights the need for prompting refinements and post-processing techniques to reduce test smells and improve long-term reliability.
@@ 7.2 Implications for Developers
Our study highlights LLM-generated test suites’ practical value when readability and maintainability are prioritized over coverage. Unlike EvoSuite, which maximizes coverage at the expense of readability, LLM-generated tests offer better structure, clearer naming, and improved comprehensibility—especially when using reasoning-based prompting like GToT and CoT. This makes these tests more accessible for long-term maintenance.
However, developers must remain aware of reliability concerns. While LLMs generate tests rapidly, compilability doesn’t guarantee execution success. Our analysis reveals many LLM-generated tests fail or time out, reinforcing the need for manual validation before pipeline integration. Although these tests achieve moderate coverage, mutation testing suggests many don’t meaningfully contribute to fault detection, often targeting trivial code portions. Developers should supplement LLM-based generation with search-based techniques to enhance fault detection.
Persistent test smells—redundant assertions, flaky tests, improper assertions—emphasize human oversight’s importance. LLMs provide a strong starting point but should be treated as assistive tools rather than autonomous solutions. Effective prompt refinement, automated filtering, and post-processing validation maximize LLM-generated tests’ utility, ensuring meaningful contributions to software quality and long-term maintainability.
@@ 8 Conclusion
This study evaluates LLMs for automated unit test generation, examining structured prompting strategies across test quality, execution reliability, maintainability, and fault detection. We analyze four LLMs and five prompting techniques to assess how prompt engineering influences test suite quality. Our findings show reasoning-based prompting, particularly GToT, improves test structure, readability, and compilability. However, execution reliability remains challenging—many tests compile but fail to verify program behavior. Mutation testing reveals tests often target trivial paths, limiting fault detection effectiveness. Hallucination-driven errors persist despite structured reasoning. While reasoning-based prompting enhances maintainability, test smells like Magic Number Tests and Assertion Roulette remain common. Coverage analysis demonstrates LLMs achieve moderate coverage but underperform EvoSuite. These findings highlight the need for hybrid approaches integrating LLM generation with search-based testing and human oversight. Our work extends beyond static correctness by incorporating execution-based classification, hallucination analysis, and mutation testing across diverse datasets. While LLMs improve test readability and structure, they still suffer from execution failures, hallucinations, and weak fault detection. Future research should explore hybrid methodologies combining LLM generation with search-based techniques, automated validation, and hallucination mitigation to enhance software quality assurance.

### v4

@@ 6.1 LLM-Based Test Generation: Strengths and Limitations
Recent studies have assessed LLMs for unit test generation along multiple dimensions, not limited to static properties. For instance, Siddiq et al. (2024a) evaluate syntactic and compilation correctness, coverage, and the presence of test smells across HumanEval and SF110, showing that while coverage can be competitive in some cases, validity and smell-related issues remain frequent. Yuan et al. (2024) conduct both a quantitative analysis and a user study, reporting that although many LLM-generated tests fail to compile or execute, the passing ones can achieve coverage and readability close to human-written tests; moreover, iterative refinement (ChatTester) improves compilability and assertion correctness. Comparative works have also positioned LLMs against SBST tools: Tang et al. (2024) examine correctness, readability, coverage, and bug detection versus EvoSuite; Yang et al. (2024) provide a broad empirical evaluation across different LLMs and datasets, concluding that LLMs still underperform EvoSuite in coverage, largely due to validity and dependency resolution issues.
Overall, prior work has already examined runtime behavior (e.g., execution and coverage) and aspects of maintainability (e.g., readability, smells). Our study complements this body of evidence by broadening scale and scope, covering three datasets with class-level generation, multiple LLM families, and five prompting strategies and by structuring failures into a fine-grained taxonomy that highlights hallucination-driven issues. We explicitly avoid exclusive claims and instead position our results as an extension and consolidation across settings and prompts.
6.2 The Impact of Prompt Engineering in Test Generation
The impact of prompt engineering on LLM-based test generation has been explored in recent research. Yuan et al. (2024) and Chen et al. (2024) examined the effects of different prompting techniques on test correctness, demonstrating that Few-Shot Learning (FSL) improves correctness but does not fully mitigate hallucinated dependencies. Similarly, Yang et al. (2024) introduced Chain-of-Thought (CoT) prompting, highlighting its ability to enhance logical reasoning in test generation.
Building upon these insights, our study systematically compares five prompting strategies namely ZSL, FSL, CoT, ToT, and GToT across multiple LLMs and datasets, focusing on structural correctness, compilability, coverage, and maintainability to assess their practical effectiveness in automated unit test generation.
@@ 6.3 Readability, Maintainability, and Test Smell Analysis
Aspects of maintainability have been discussed in prior work, though often in limited scope. Siddiq et al. (2024a) quantify the prevalence of smells such as duplicated asserts and empty tests in LLM-generated suites, reporting nontrivial rates particularly on SF110. Yuan et al. (2024) complement this with a user study indicating that passing LLM-generated tests can achieve readability close to human-written ones, while also noting frequent failures and incorrect assertions in the broader population. Beyond Java, Schäfer et al. (2023) show that LLM-generated JavaScript tests can include meaningful oracles and competitive coverage, though such results depend strongly on runtime and language-specific contexts.
Building on these observations, our study provides a broader and more systematic assessment of maintainability. We analyze test smells using TsDetect, evaluate readability with a specialized deep learning model for JUnit tests (Scalabrino et al., 2018), and incorporate structural measures such as Match Success Rate (MSR) and Code Extraction Success Rate (CSR). In contrast to prior evaluations, we scale this analysis to three datasets with class-level test generation, multiple LLM families, and five prompting strategies, explicitly examining how reasoning-based prompting modulates maintainability indicators. This comprehensive approach uncovers structural and readability challenges in LLM-generated tests and provides actionable insights for prompt design and post-processing aimed at improving long-term usability.
@@ 6.4 Advancing LLM-Based Unit Test Improvement
Recent studies have explored LLM-driven unit test improvement techniques, primarily focusing on iterative refinement. Gu et al. (2024) introduced TestArt, a co-evolutionary approach that iteratively refines LLM-generated tests through repair iterations, enhancing correctness over multiple steps. Similarly, Zhang et al. (2025) proposed CITYWALK, which employs iterative refinement guided by contextual cues to improve generated tests, with a focus on increasing functional correctness. Alshahwan et al. (2024) investigated LLM-driven test improvements at Meta, focusing on refining correctness but without addressing key aspects such as maintainability, readability, or test smells. These studies demonstrate progress in test improvement but remain limited to correctness-focused refinements.
In contrast, our work provides a multi-dimensional assessment that goes beyond correctness-focused refinements. We complement prior improvement techniques by evaluating structure- and maintenance-oriented qualities: extractability (MSR/CSR), syntactic correctness, compilability, coverage, readability, and test smells. Moreover, we introduce readability and maintainability metrics to assess whether generated tests are understandable and practically usable in real-world software projects.
Our findings indicate that while iterative refinement can enhance the syntactic and structural quality of LLM-generated tests, it does not fully address maintainability and structural robustness: compilability remains fragile and test smells persist across datasets and prompting strategies. This suggests the need for hybrid approaches that combine Search-Based Software Testing (SBST) with LLM-based generation to jointly target correctness and maintainability.
@@ 6.5 Key Takeaways and Research Gap
Prior works on LLM-based test generation primarily focus on syntactic correctness and code coverage, leaving key aspects of readability, maintainability, and structural reliability underexplored. Our study fills this gap by providing a comprehensive evaluation framework that goes beyond static correctness checks to assess multiple dimensions of test quality. Specifically, we evaluate test readability and maintainability using deep learning-based readability scoring, aligned with human evaluation guidelines.
Unlike previous studies that explore only one or two prompting techniques, we systematically compare five strategies—ZSL, FSL, CoT, ToT, and GToT. We also integrate mutation testing, a missing component in prior research, to evaluate fault detection capabilities. Beyond correctness and structural coverage (line, instruction, method), we analyze test smells, code extraction reliability (MSR & CSR), compilation errors, and runtime failures, uncovering key challenges in dependency resolution and API usage. Unlike previous studies that explore only one or two prompting techniques, we systematically compare five strategies(ZSL, FSL, CoT, ToT, and GToT) across four LLMs and three datasets. Beyond structural correctness and coverage (line, instruction, and method), we analyze code extractability (MSR & CSR), compilation success, and test smells to uncover recurring issues in dependency resolution and API usage.
A key differentiator of our work is the diversity of datasets: SF110 (academic), Defects4J (real-world bugs), and CMD (industrial codebase). This design ensures that our findings generalize across heterogeneous software environments. By bridging structural reliability, readability, maintainability, and dataset diversity, our study establishes a broad, data-driven foundation for understanding the practical effectiveness and current limitations of LLM-generated test suites.
@@ 7 Discussion of Findings and Implications
Our study reveals that reasoning-based prompting (CoT, GToT) improves the structural correctness and maintainability of LLM-generated tests, yet reliability and integration robustness remain major challenges. While many tests compile successfully, they often exhibit dependency mismatches or redundant logic, and test smells continue to affect maintainability. Hybrid strategies combining Search-Based Software Testing (SBST) with LLM-driven generation and targeted human oversight appear essential for achieving higher robustness. Although prompt engineering enhances structural reliability, hallucination-induced errors persist, highlighting the need for stronger contextual grounding and improved handling of dependencies and encapsulation. These findings underscore the importance of adaptive test generation pipelines that integrate automated validation, dependency resolution, and iterative refinement to improve the practical usability of LLM-generated unit tests.
@@ 7.1 Implications for Researchers
Our findings emphasize reasoning-based prompting’s significant role in improving LLM-generated test suites. Unlike prior studies (Siddiq et al., 2024a; Tang et al., 2024; Yang et al., 2024), which did not extensively analyze prompt engineering, we systematically evaluate ZSL, FSL, CoT, ToT, and GToT. Results show that structured reasoning significantly enhances test adherence, compilability, and structural reliability, with GToT consistently outperforming other strategies.
Our multi-level coverage analysis reveals that LLM-generated tests achieve moderate coverage but remain inferior to EvoSuite, reinforcing the need for hybrid approaches that integrate SBST techniques with LLM-based generation. Our findings also challenge the common assumption that larger models inherently produce better tests: GPT-3.5-turbo occasionally outperforms GPT-4 on CMD, highlighting the need to better understand how architectural and dataset characteristics influence performance.
Beyond correctness and coverage, our readability and maintainability analysis reveals that while LLMs produce syntactically correct and readable tests, readability alone does not guarantee maintainability. Test smells—particularly Magic Number Tests and Assertion Roulette—persist across models and prompting strategies, increasing debugging complexity. These results highlight the need for more refined prompting strategies and automated post-processing techniques to reduce test smells and improve long-term test reliability.
@@ 7.2 Implications for Developers
Our study highlights the practical value of LLM-generated test suites when readability and maintainability are prioritized over raw coverage. Unlike EvoSuite, which maximizes coverage at the expense of readability, LLM-generated tests generally exhibit clearer structure, more meaningful naming, and higher comprehensibility—particularly when reasoning-based prompting techniques such as GToT and CoT are applied. This makes these tests more accessible for long-term maintenance and collaborative development.
However, developers must remain aware of reliability concerns. While LLMs can generate large numbers of tests rapidly, issues such as partial compilability, redundant assertions, or unverified logic can still arise. Persistent test smells—including redundant or fragile assertions, flaky behavior, and inconsistent naming—emphasize the importance of human oversight. LLMs thus serve best as “assistive tools” rather than fully autonomous test generators. Careful prompt refinement, automated filtering, and lightweight post-processing validation can maximize their practical utility, ensuring that generated tests meaningfully contribute to software quality and long-term maintainability.
@@ 8 Conclusion
This study presents a large-scale evaluation of Large Language Models (LLMs) for automated unit test generation, systematically analyzing how structured prompting strategies influence test quality, readability, and maintainability across multiple datasets. By comparing five prompting techniques (ZSL, FSL, CoT, ToT, and GToT) and four LLMs, our results show that reasoning-based prompting especially GToT and CoT substantially improves test structure, compilability, and readability compared to zero-shot prompting. However, LLM-generated tests remain limited in robustness and reliability. Although they exhibit cleaner code and higher readability than traditional SBST-generated tests, persistent issues such as incomplete imports, redundant assertions, and recurring test smells (e.g., Magic Number Tests, Assertion Roulette) reduce their long-term maintainability. Coverage analysis further reveals that while LLMs achieve moderate coverage, they still underperform EvoSuite in exploring program structure. Overall, our findings indicate that LLMs can indeed assist in automated test generation, but they are not yet ready to replace traditional testing tools. Their greatest strength lies in producing human-readable, maintainable test scaffolds that can complement existing SBST or developer-written tests. To unlock their full potential, future research should focus on hybrid methodologies that integrate LLM-based generation with automated validation, dependency resolution, and search-based refinement. Such approaches could enable LLMs to evolve from assistive code generators into reliable components of software quality assurance pipelines.